4. Solve for x: log v3 x + log v3 (x + 8) = 2
The v3 parts are supposed to be like the reverse of an exponent. They are at the bottom and not the top.

Answer by venugopalramana(3286)

4. Solve for x: log v3 x + log v3 (x + 8) = 2
The v3 parts are supposed to be like the reverse of an exponent. They are at the bottom and not the top.
USE THE FORMULAE
1.LOG(X) TO BASE B + LOG(Y) TO BASE B = LOG(X*Y) TO BASE B
2.LOG(X) TO BASE B = P MEANS B^P=X
LOG(X) TO BASE 3 +LOG(X+8) TO BASE 3 = LOG[(X)(X+8)]TO BASE 3 =2
HENCE
3^2=(X)(X+8)=X^2+8X
X^2+8X-9=0
X^2+9X-X-9=0
X(X+9)-1(X+1)=0
(X-1)(X+9)=0
HENCE X-1=0......OR....X=1...........SINCE X CANNOT BE -9 AS LOG OF A NEGATIVE NUMBER DOES NOT EXIST.
CHECK
LOG(1) TO BASE 3 =0
LOG(1+8) TO BASE 3 = 2 SINCE 3^2=9